What is the EMV for Option 1 in the following decision table?
  States of Nature  
Alternatives S1 S2
P .3 .7
Option 1 15,000 20,000
Option 2 10,000 30,000

 

A) 18,500 B) 17,000 C) 20,000 D) 17,500 E) 15,000
Answer #1

A) 18,500

For Option 1

S1 = 0.3*15000

S2 = 0.7*20000

S1 + S2 = 0.3*15000 + 0.7*20000

= 4500+14000

= 18500